It took a train 45 seconds to pass through a 1320-foot tunnel. At the same speed, it took 15 seconds for the train to pass a pole on the side of the tracks. How long was the train?
Answer by mananth(16946)

The distance the train has to travel to get out of the tunnel is = length of train + length of tunnel
let length of train be x feet
x+1320 = total distance it travels
speed = x+1320/45
when crossing a pole speed = x/15
speeds at both the times are same
x/15 = x+1320/45
45x=15(x+1320)
45x=15x+19800
30x= 19800
x= 19800/30
660 feet is the length of the train
Ananth